Mr. Gulsby had $15,000 to invest, part at 12% and the rest at 6%. If he wants an annual return of at least $1,500, what amount should he invest at each rate?
Answer by stanbon(75887)

Mr. Gulsby had $15,000 to invest, part at 12% and the rest at 6%. If he wants an annual return of at least $1,500, what amount should he invest at each rate?
--------------------------
Quantity Equation: x + y = 15000
Interest Equation:0.12x + 0.06y = <= 1500
-------------------------------------------
Multiply 1st by 12
Multiply 2n by 100
-----------------------
12x + 12y = 12*15000
12x + 6y <= 10*15000
===================
Subtract 2nd from 1st and solve for "y":
6y = 2*15000
y = 5000 (amount invested at 6%
Since x + y = 15000, x = 10,000 (amount to be invested at 12%)
====================================================----------
Cheers,
Stan H.